OpenStudy (anonymous):

(1/9)l5x-3l -3 ≥ 2 (Those l's are for absolute value. I am so confused. Anyone got a solution?)

OpenStudy (mathmate):

Solve separately the expression |5x-3| as +(5x-3) and -(5x-3), which means you have two solutions. solve (1/9)(5x-3)\(\ge\)2 and -(1/9)(5x-3)\(\ge\)2 to get the two required solutions.

OpenStudy (anonymous):

I have to solve for x and I think I have to start with l5x-3l - 3 = 18

OpenStudy (mathmate):

Yes, that's ok too!, now put the expression as 5x-3 and -(5x-3) and solve the equations.

OpenStudy (mathmate):

but (1/9)(5x-3)-3≥2 => (5x-3)-27≥18 and -(1/9)(5x-3)-3≥2 => -(5x-3)-27≥18

OpenStudy (anonymous):

ummm

OpenStudy (mathmate):

Add 27 to both sides, we have (5x-3)-27+27≥18+27 -(5x-3)-27+27≥18+27 or 5x-3≥45 and -5x+3≥45 solve accordingly.